You are required to specify the power rating of a resistor, R1, to be used in a car engine management system, so that an order can be placed for the component (resistor).

The resistance required is 2Ω. The maximum current that will be flowing in the resistor will be 10A.

This resistor is only available in the following power ratings - Choose the most appropriate power rating for this resistor for the given application.


1. 2W

2. 20W

3. 120W

4. 250W
1
Expert's answer
2015-01-12T07:37:01-0500
Solution:
Givem:
I=10A
R=2Ohm

P=R*I^2 = 2*10^2 = 200 W
So in our case 120W is best variant
Answer 3